What is the bond equivalent yield of a bond if it has 200 days to maturity, a par value of $10000, and is currently trading at $9780? Enter your answer as a decimal, rounded to 4 decimal places. Your Answer: Answer Question 4 (1 point) suppose that short-term municipal bonds currently offer yields of 4%, while comparable taxable bonds pay 5%, which gives you the higher after-tax yield if your tax bracket is 30%? Municipal bonds Taxable bonds Question 5 (1 point) The current yield on a municipal bond is 5.5% . What would be the equivalent taxable yield of this bond to a taxpayer in a 35% tax bracket? Enter your answer as a decimal, not percentage (e.g., 0.05 instead of 5%), and round your answer to 4 decimal places.
Answer #1

Question 3 :

Par value at redemption = $ 10000

bond price = $ 9780

Number of days of maturity = 200 days

Profit of holding this bond = Par value at redemption - Bond purchase price

= $ 10,000 - $ 9,780

= $ 220

Yield from the bond for 200 days = profit from holding the bond / Purchase price of the bond

= $ 220 / $ 9780 * 100

= 0.0224949 * 100

   = 2.2495

Yield annualized = 2.2495 * 365 days / 200 days

= 4.1053 % or ( 0.0410)

question 4 :

Yield from municipal bond = 4 % ( tax free yield bond)

Yield from taxable bond before tax = 5 %

Yield from taxable bond after tax = 5 % * ( 1 - tax rate )

= 5 * ( 1- 30%)

= 5% * ( 0.70)

= 3.5 %

Yield from municipal bond = 4 %

Yield from taxable bond = 3.5 %

Yield from municipal bond is more than yield from taxable bond. ( 4% > 3.5%)

option 1

Question 5 :

Equivalent taxable yield of bond = Yield from municipal bond / (1- tax rate)

= 5.5 % / ( 1 - 35)

= 5.5 % / ( 65)

= 5.5 % / (65)

= 0.0846 ( 8.46 % )